the line through ( -2,5)with slope -1
Answer by nerdybill(7384)

the line through ( -2,5)with slope -1
.
Plug given data into "point-slope" form:
y - y1 = m(x - x1)
y - 5 = -1(x - (-2))
y - 5 = -(x + 2)
y - 5 = -x - 2
y = -x + 3 (this is what they're looking for)